OpenStudy (anonymous):

cal 1 problem -.5<(1/x)-(1/5)<.5 -.3<1/x<.7 -3.33

OpenStudy (amoodarya):

$$ -.5<\frac{1}{x}-\frac{1}{5}<.5\\ \frac{-1}{2}<\frac{1}{x}-\frac{1}{5}<\frac{1}{2}\\ \frac{-1}{2}+\frac{1}{5}<\frac{1}{x}<\frac{1}{2}+\frac{1}{5}\\ \frac{-3}{10}<\frac{1}{x}<\frac{7}{10}\\ \frac{1}{x}<\frac{7}{10}\rightarrow \frac{1}{x}-\frac{7}{10}<0 \\ \frac{-3}{10}<\frac{1}{x}\rightarrow 0<\frac{1}{x}+\frac{3}{10}\\ $$